OpenStudy (anonymous):

compute the division 14/ 4\5

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

if you mean 14/ (4/5) then it is 17.5

OpenStudy (anonymous):

don't

jhonyy9 (jhonyy9):

14/4/5 = (14/4 )x(1/5) =14/20 =7/10